# Inferring Pluggable Types with Machine Learning

### v1

@@ V NaP-ASTs for Nullability
With the dataset in section IV, we trained the graph-based models (GCN and GTN) by following the NaP-AST encoding method described in section III-A. Our code, dataset, and experimental scripts are open-source and available [23]. We call the resulting models NullGCN and NullGTN.
V-A NaP-AST Pruning
@@
In this section, we describe the specific pruning strategies that we used for each pruning phase described in section III-A.
V-A1 Phase 1: Pruning by Proof
@@
Since primitive types cannot store null values in Java, we exclude them from the nodes that the model is trained on. Nullability is relevant to most of the program, but we expect that for other pluggable type systems that target rarer problems, this step could prune dramatically-larger parts of the program. For example, an instantiation of our methodology for a type system that detects resource leaks could remove all parts of the program that definitely do not manipulate resources (i.e., most of the program!).
V-A2 Phase 2: Pruning by Node Type
First, we conducted an ablation study over node types. We dropped each node type and calculated the F1 score of a preliminary model on held-out testing data.The results are plotted in fig. 4 (test runs were repeated 50 times each to control for the statistical nature of the models). Note “<random>” in red in fig. 4: this represents dropping nodes at random from the graph, without regard to their type. We discarded all nodes of the types that appear to the left of “<random>” in fig. 4: that is, node types whose removal damages the model’s predictive ability less than removing random nodes. We connect the children of dropped nodes to their parents.
For example, since humans often write @Nullable annotations on method declaration nodes (to indicate nullable return types), dropping them dramatically reduces the F1 score to about 0.66 . By contrast, node types like MemberValuePair and IntersectionType seem to have no measurable impact on the performance of the model, which makes sense: value pairs are unrelated to reasoning about nullness. As another example, IntersectionType nodes indicate a generic type construct—which is entirely orthogonal to nullness.
Dropping some node types suprisingly drops the F1 score (if ClassOrInterfaceDeclaration is dropped, for example, the F1 score falls to 0.76). We determined via manual investigation that nodes of this type help group the methods of nested classes, so even though they are rarely or never nullable themselves, retaining them is important to model performance. Another node type that helps group nodes together is BlockStmt; dropping them reduces the F1 score to 0.77.
InstanceOfExprArrayInitializerExprTypeExprMethodCallExprForStmtVariableDeclaratorUnaryExprVarTypeNullLiteralExprNormalAnnotationExprIfStmtLambdaExprArrayCreationLevelEnclosedExprBinaryExprImportDeclarationExplCtorInvocStmtReturnStmtTryStmtForEachStmtIntegerLiteralExprWildcardTypeConstructorDeclarationLineCommentConditionalExprObjectCreationExprArrayType<random>ArrayCreationExprArrayAccessExprCharLiteralExprIntersectionTypeAssignExprSingleMemAnnoExprMemberValuePairClassExprParameterEnumDeclarationBlockStmtTypeParameterClassOrInterfaceDeclModifierThisExprMethodDeclaration0.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9Node TypeF1 after dropping
@@ V-A2 Phase 2: Pruning by Node Type
Figure 4: Ablation study to decide which node types can be dropped without affecting the F1 score. Only node types that performed worse than dropping random nodes (in red) were retained in later models.
V-A3 Phase 3: Pruning By Subtree
@@
Using a preliminary model, we conducted an ablation study to determine which kinds of statement nodes we can remove, if the corresponding subtree does not contain a NullLiteralExpr node. In the ablation study, we pruned all such subtrees under a particular kind of statement node (and all nodes connected to that subtree via a name layer), for each kind of statement. The resulting F1 scores for the preliminary model appear in fig. 5. Our NaP-AST construction procedure prunes the subtrees and name-connected nodes for the types of statement nodes that appear to the left of the “<random>” entry (in red) in fig. 5: that is, those statement types that are less useful to the model than a random statement. For example, a nullability NaP-AST will not contain a for loop (ForStmt in fig. 5) nor nodes for any of the variables manipulated in that loop unless there is a null literal within the loop. However, all expression statements (not within the subtree of another pruned statement) will be preserved, whether or not the ExpressionStmt node for the statement has a null literal in one of its subtrees—removing expression statements damages model performace too much.
ForStmtBlockStmtLocalClassDeclStmtSynchronizedStmtForEachStmtCatchClauseWhileStmtLabeledStmtAssertStmtThrowStmt<random>ExplCtorInvocStmtReturnStmtIfStmtContinueStmtSwitchStmtTryStmtExpressionStmt0.80.80.80.80.850.850.850.850.90.90.90.90.950.950.950.95Node TypeF1 after dropping
@@
Figure 5: Results of the ablation study described in section III-A5 by statement type. Subtrees of nodes of the types to the left of “<random>”, and nodes connected to those subtrees by the name layer, are pruned from the NaP-AST.
TABLE I: NullAway benchmark results. “War.” means “Warnings.” “Red. %” means “reduction percentage”: that is, the reduction in warnings achieved by NullGTN vs. the unannotated code.
NullGCN: Magicoder: NullGTN:
Project War. Red. % Precision Recall War. Red. % Precision Recall War. Red. % Precision Recall
ColdSnap 2 99% 4 8 Broken NA NA NA 2 99% 15 90
QRContact 0 - 0 0 Broken NA NA NA 0 - 11 61
caffeine 117 -11% 83 44 6 94% 0 0 73 30% 85 95
jib 122 -20% 84 19 Broken NA NA NA 101 0% 62 91
keyvaluestore 44 -450% 13 70 Broken NA NA NA 4 50% 22 100
meal-planner 11 -1000% 0 0 Broken NA NA NA 1 0% 10 100
uLeak 6 50% 37 100 12 0% 0 0 8 33% 75 100
AutoDispose 30 -1400% 3 11 Broken NA 0 0 2 0% 0 0
picasso 0 - 81 44 0 - 25 1 0 - 0 0
RIBs 0 100% 100 100 1 0% 0 0 1 0% 0 0
skaffold-tools-for-java 1 0% 100 83 1 0% 0 0 1 0% 0 0
butterknife 4 0% 100 50 4 0% 100 50 0 100% 100 100
Totals 337 31% .51 .32 NA NA .2 .03 193 60% .6 .89
V-A4 Post-processing
@@
We added four nullability-specific post-processing steps to ensure internal consistency.
@@
Inheritance
A method’s inferred parameter or return types in the same inheritance hierarchy may not match. We remove @Nullable annotations in superclasses/interfaces no such annotation was inferred in the subclass.
V-B Training
TABLE II: GTN hyperparameters
@@ V-B Training
We trained the GTN on a cluster with Intel Xeon E5-2630-v4 CPUs (2.2 GHz), 2 GPUs and 50 GB RAM. Since the adjacency matrices that we use to represent the AST are sparse, they take up a lot of space on the GPU. So, we were only able to train a sample consisting of 1.15% of our dataset at a time. We thus chose batch training, which allows us to train the GTN on the full dataset over a longer period of time. We fixed the input size to 8000 nodes, and only trained on as many classes at a time as fit into that space. The GCN’s training (and all the experiments in section VI) were run on an AWS g4dn.xlarge instance and a Deep Learning PyTorch AMI.
V-B1 Hyperparameters
@@
After some trial and error, we used a simple 2 layer GCN with batching. The GTN’s trained hyperparameters are in table II. These hyperparameters were chosen by manual grid search. For this dataset, non-local weight of -2 performs well independent of other variables. Since we trained a large dataset in batches, we chose a low initial learning rate and used weight decay throughout the training process to avoid overfitting.
V-C Thresholding during Deployment
@@
The models are trained on an equal number of nodes that are nullable and non-null, and all of its training data was sourced from classes that contained at least one nullable member. Model predictions reflect this: averaged over many pieces of code, the model’s average nullability prediction is about half nullable elements and half non-null elements. In real code, however, there are many more non-null elements than nullable elements: in fact, real code contains many classes that have no nullable members at all! To account for this, when applying each model, we use a thresholding strategy: we only actually mark elements as nullable (by writing an @Nullable annotation) when the model is very confident in that element’s nullability. In our experiments with real projects in section VI-A, we applied a threshold of 90% confidence. In a deployment of one of these models on a previously-unannotated codebase (which is our envisioned deployment scenario), we recommend using a similar threshold.
VI Experiments and Results
We address the research questions from section I:
  • RQ1: What kind of ML model should we use, and how should we post-process its output?
  • RQ2: How should we encode the program in a type-system-specific way to train the model?
@@ VI Experiments and Results
  • RQ3: How much data do we need to make that model effective at inferring qualifiers in practice?
@@
We performed two experiments. The first is a large-scale evaluation of each model’s ability to infer ground-truth, human-written annotations on real open-source projects from a previous study of a nullability checker (section VI-A) to answer RQ1 and RQ2. The second experiment (section VI-B) evaluated how much training data the best model from section VI-A required for good performance, to answer RQ3.
VI-A NullAway Benchmark
@@
We ran all three models on the open-source benchmark the paper about the NullAway pluggable typechecker [2]. This benchmark (from their artifact [24]) is a collection of 18 projects from GitHub in 2019; each was annotated with nullability annotations when the artifact was published. The artifact also includes an evaluation tool to run NullAway on these projects. Two of these repositories no longer exist;we could build the project and run NullAway on 14 of the remaining 16; two more projects no longer use @Nullable annotations (in the branches pulled by the artifact’s evaluation tool).We consider the remaining 12 projects that we can typecheck using NullAway; however, we note that of these projects, 3 of the 12 do not typecheck with no warnings, even with the human-written annotations. We assume that this is because of either changes in the projects or changes in NullAway itself.
@@
We checked each benchmark project with NullAway in three configurations: 1) with the human-written annotations that came with the projects, 2) after removing the human-written annotations from each project, and 3) after using each model to re-annotate the code. We measured two proxies for human annotation effort: each model’s precision/recall with respect to the human-written annotations (which are erased before passing the projects to each model) and the reduction in the warnings that NullAway issues using our annotations versus the unannotated code. The results appear in table I.
VI-A1 NullGCN
@@
We trained NullGCN using PyTorch and the PyTorch Geometric library. It uses two convolutional layers, dropout regularization, negative log-likelihood loss function and the Adam optimizer. It achieves an accuracy of 0.6 on the test set during training. On the full NullAway benchmark, NullGCN’s performance is not particularly good: its recall is 32%, precision is 51%, and warning reduction is only 31%. While this is better than no inference at all, this result suggests to us that the hypergraph structure of the GTN (which does much better) is important to modeling where humans write annotations: the same graph is not as useful if the edges are not known to be of different types.
VI-A2 Magicoder LLM
We use Magicoder to reannotate the NullAway benchmark projects. Magicoder uses text for input and output, we did not use the NaP-AST encoding.Magicoder seems to perform well on one project (caffeine). However, it often breaks the code in unpredictable ways by making seemingly-nonsensical changes. Even when this does not occur, the LLM’s token limit is too restrictive for it to be effective on most large projects. These problems together lead to its low precision (20%) and recall (3%). While Magicoder appears to do well on the seemingly-difficult caffeine project in table I, the results are difficult to credit: while the resulting code builds with NullAway, and NullAway issues only 6 warnings on it (a seeming warning reduction of 94%!), the structure of the code was altered too much to tell which node goes where: that is, we cannot even compute precision and recall, because Magicoder’s edits to caffeine are too extensive. The LLM did not follow the prompt to not change the code except for placing annotations.
VI-A3 NullGTN
@@
We built our GTN model (NullGTN) by modifying a fork of the original code [25]. It takes an average of about 16 hours to train the model. The largest input source code we evaluated (6,470 non-comment, non-blank lines) took about 5 minutes for inference. NullGTN achieved an F1 score of .95 on the test set during training. On the NullAway Benchmark, NullGTN discovered 89% of the human-written type qualifiers—most of them, with a precision of 60%. However, this only eliminated 60% of the NullAway warnings that a human would need to resolve to adapt these benchmarks to use NullAway—a large amount, but far from all of the warnings. One source of NullGTN’s inability to reduce warnings further is imprecision, which leads to extra annotations, which leads to extra warnings. Another cause of low warning reduction for NullGTN is interactions with libraries. NullGTN has no model of what libraries do, and unlike an LLM it cannot guess based on e.g., the names of library methods. Both the jib and caffeine projects (where NullGTN fails to remove many warnings despite good recall) use libraries extensively: caffeine heavily uses Google Guava, while jib is a modular system with many APIs and plugins, which the model was unaware of.
Overall, these results are encouraging: NullGTN makes a good start on bringing these benchmarks into compliance with NullAway, and finds the vast majority (89%) of the annotations that the developers themselves keep in their codebases. Our answer to RQ2 is therefore positive: NullGTN can make significant progress towards annotating a codebase for a typechecker, and many of the annotations that it produces are remarkably similar to those a human developer would write.
102030405060708090100200200200200400400400400600600600600Training percentageWarnings
Figure 6: Number of warnings by percentage of data trained.
VI-B RQ3: How much data do we need to make that model effective at inferring qualifiers in practice?

### v2

@@ 5. NaP-ASTs for Nullability
With the dataset in section 4, we trained graph-based models (GCN and GTN) using our NaP-AST encoding method (section 3.2). This section describes the NaP-AST construction process; we call the resulting models NullGCN and NullGTN.
5.1. NaP-AST Pruning
@@
In this section, we describe the specific pruning strategies that we used for each pruning phase described in section 3.2.
5.1.1. Phase 1: Pruning by Proof
@@
Since primitive types cannot store null values in Java, we exclude them from the nodes that the model is trained on. Nullability is relevant to most of the program, but we expect that for other pluggable type systems that target rarer problems, this step could prune dramatically-larger parts of the program. For example, an instantiation of our methodology for a type system that detects resource leaks could remove all parts of the program that definitely do not manipulate resources (i.e., most of the program).
5.1.2. Phase 2: Pruning by Node Type
InstanceOfExprArrayInitializerExprTypeExprMethodCallExprForStmtVariableDeclaratorUnaryExprVarTypeNullLiteralExprNormalAnnotationExprIfStmtLambdaExprArrayCreationLevelEnclosedExprBinaryExprImportDeclarationExplCtorInvocStmtReturnStmtTryStmtForEachStmtIntegerLiteralExprWildcardTypeConstructorDeclarationLineCommentConditionalExprObjectCreationExprArrayType¡random¿ArrayCreationExprArrayAccessExprCharLiteralExprIntersectionTypeAssignExprSingleMemAnnoExprMemberValuePairClassExprParameterEnumDeclarationBlockStmtTypeParameterClassOrInterfaceDeclModifierThisExprMethodDeclaration0.60.60.70.70.80.80.90.9Node TypeF1 after dropping
@@ 5.1.2. Phase 2: Pruning by Node Type
Figure 5. Ablation study to decide which node types can be dropped without affecting the F1 score. Only node types that performed worse than dropping random nodes (in red) were retained in later models.
First, we conducted an ablation study over node types. We dropped each node type and calculated the F1 score of a preliminary model on held-out testing data.The results are plotted in fig. 5 (test runs were repeated 50 times each to control for the statistical nature of the models). Note “¡random¿” in red in fig. 5: this represents dropping nodes at random from the graph, without regard to their type. We discarded all nodes of the types that appear to the left of “¡random¿” in fig. 5: that is, node types whose removal damages the model’s predictive ability less than removing random nodes. We connect the children of dropped nodes to their parents.
For example, since humans often write @Nullable annotations on method declarations (to indicate nullable return types), dropping them dramatically reduces the F1 score to about 0.66 . By contrast, node types like MemberValuePair and IntersectionType seem to have no measurable impact on the performance of the model, which makes sense: value pairs are unrelated to reasoning about nullness. As another example, IntersectionType nodes indicate a generic type construct—which is entirely orthogonal to nullness.
Dropping some node types surprisingly drops the F1 score (if ClassOrInterfaceDeclaration is dropped, for example, the F1 score falls to 0.76). We determined via manual investigation that nodes of this type help group the methods of nested classes, so even though they are rarely or never nullable themselves, retaining them is important to model performance. Another node type that helps group nodes together is BlockStmt; dropping them reduces the F1 score to 0.77.
5.1.3. Phase 3: Pruning By Subtree
@@
Using a preliminary model, we conducted an ablation study to determine which kinds of statement nodes we can remove, if the corresponding subtree does not contain a NullLiteralExpr node. In the ablation study, we pruned all such subtrees under a particular kind of statement node (and all nodes connected to that subtree via a name layer), for each kind of statement. The resulting F1 scores for the preliminary model appear in fig. 6. Our NaP-AST construction procedure prunes the subtrees and name-connected nodes for the types of statement nodes that appear to the left of the “¡random¿” entry (in red) in fig. 6: that is, those statement types that are less useful to the model than a random statement. For example, a nullability NaP-AST will not contain a for loop (ForStmt in fig. 6) nor nodes for any of the variables manipulated in that loop unless there is a null literal within the loop. However, all expression statements (not within the subtree of another pruned statement) will be preserved, whether or not the ExpressionStmt node for the statement has a null literal in one of its subtrees—removing expression statements damages model performance too much.
ForStmtBlockStmtLocalClassDeclStmtSynchronizedStmtForEachStmtCatchClauseWhileStmtLabeledStmtAssertStmtThrowStmt¡random¿ExplCtorInvocStmtReturnStmtIfStmtContinueStmtSwitchStmtTryStmtExpressionStmt0.80.80.850.850.90.90.950.95Node TypeF1 after dropping
@@
Figure 6. Results of the ablation study in section 3.2.5 by statement type. Subtrees of nodes of types to the left of “¡random¿”, and nodes connected to those subtrees by the name layer, are pruned from the NaP-AST.
5.1.4. Post-processing
@@
Four nullability-specific post-processing steps ensure internal consistency.
@@
Table 1. GTN hyperparameters
@@
Inheritance
A method’s inferred parameter or return types in the same inheritance hierarchy may not match. We remove @Nullable annotations in superclasses and interfaces when no such annotation was inferred in the subclass or implementing class, respectively.
6. Training
@@
We trained the GTN on a cluster with Intel Xeon E5-2630-v4 CPUs (2.2 GHz), 2 GPUs and 50 GB RAM in the Lochness HPC. Since the adjacency matrices that we use to represent the AST are sparse, they take up a lot of space on the GPU. So, we were only able to train a sample consisting of 1.15% of our dataset at a time. We thus chose batch training, which allows us to train the GTN on the full dataset over a longer period of time. We fixed the input size to 8000 nodes, and only trained on as many classes at a time as fit into that space. The GCN’s training (and all the experiments in section 7) were run on an AWS g4dn.xlarge instance and a Deep Learning PyTorch AMI. We trained NullGCN using PyTorch and the PyTorch Geometric library. The GCN uses two convolutional layers, dropout regularization, negative log-likelihood loss function and the Adam optimizer.
6.0.1. Hyperparameters
@@
After some trial and error, we used a simple 2 layer GCN with batching. The GTN’s trained hyperparameters are in table 1. These hyperparameters were chosen by manual grid search. For this dataset, non-local weight of -2 performs well independent of other variables. Since we trained a large dataset in batches, we chose a low initial learning rate and used weight decay throughout the training process to avoid overfitting.
6.1. Thresholding during Deployment
@@
The models are trained on an equal number of nodes that are nullable and non-null, and all the training data was sourced from classes that contained at least one nullable member. Their predictions reflect this: averaged over many pieces of code, the average nullability prediction is about half nullable elements and half non-null elements. In real code, however, there are many more non-null elements than nullable elements: in fact, real code contains many classes that have no nullable members at all! To account for this, when applying each model, we use a thresholding strategy: we only actually annotate elements as nullable when the model is very confident in that element’s nullability. In our experiments with real projects in section 7.1, we applied a threshold of 90% confidence. In a deployment of one of these models on a previously-unannotated codebase (which is our envisioned deployment scenario), we recommend using a similar threshold.
7. Experiments and Results
We address the research questions from section 1:
  1. RQ1
    What kind of machine-learning model architecture is suitable for inferring pluggable types, and how does that model differ from those in prior type inference works for other languages?
  2. RQ2
@@
    How much data is needed to make that model effective at inferring qualifiers in practice?
@@
We performed two experiments. The first is a large-scale evaluation of each model’s ability to infer ground-truth, human-written annotations on real open-source projects from a previous study of a nullability checker (section 7.1) to answer RQ1. The second experiment (section 7.3) evaluated how much training data the best model from section 7.1 required for good performance, to answer RQ2. All of our code, dataset, and experimental scripts are open-source and available (Kazi Siddiqui, Martin Kellogg, [n. d.]).
7.1. NullAway Benchmark Experiment
@@
We ran each model on the open-source benchmark from the NullAway paper (Banerjee et al., 2019b). This benchmark (from their artifact (Banerjee et al., 2019a)) is a collection of 18 projects from GitHub; each was annotated with nullability annotations when the artifact was published. The artifact also includes an evaluation tool to run NullAway on these projects. Two of these repositories no longer exist; we could build the project and run NullAway on 11 of the remaining 16; two more projects no longer use @Nullable annotations (in the branches pulled by the artifact’s evaluation tool). We consider the remaining 9 projects that we can typecheck using NullAway; however, we note that of these projects, 2 of the 9 do not typecheck with no warnings, even with the human-written annotations. We assume that this is because of either changes in the projects or changes in NullAway itself.
In the remaining projects, 142 out of the 594 classes (~23%) have a @Nullable annotation. This is a typical ratio; as another point of comparison, consider the eureka and spring-framework projects (both of which are outside the dataset but are annotated for NullAway). Eureka  [^5] https://github.com/Netflix/eureka has 21/408 annotated classes (~5%); spring-framework  [^6] https://github.com/spring-projects/spring-framework has 3016/8800 (~34%). These projects are examples of human-annotated projects near the extremes of the distribution.
@@
We checked each benchmark project with NullAway in three configurations: 1) with the human-written annotations that came with the projects, 2) after removing the human-written annotations from each project, and 3) after using each model to re-annotate the code. We measured two proxies for human annotation effort: each model’s precision/recall of the human-written annotations (which are erased before passing the projects to each model) and the reduction in the warnings from NullAway using our annotations versus the unannotated code. The results appear in table 2.
Table 2. NullAway benchmark results. “War.” means “Warnings.” “Red. %” means “reduction percentage”: that is, the reduction in warnings achieved by each model vs. the unannotated code.
Unannotated NullGCN TDG GCN NullGTN GPT-4 CodeQwen
Project War. War. Red. % Prec. Rec. War. Red. % Prec. Rec. War. Red. % Prec. Rec. War. Red. % Prec. Rec. War. Red. % Prec. Rec.
ColdSnap 256 2 99% .06 .2 256 0% .72 .28 2 99% .15 .8 0 100% .65 .28 2 99% 1 .26
QRContact 0 0 - 0 0 0 0% .66 1 0 - 0 0 0 - 1 1 0 - 0 0
jib 101 122 -20% .62 .24 101 0% .9 .33 101 0% .55 .68 166 -64% .82 .33 101 0% .9 .28
meal-planner 1 11 -1000% .07 .33 1 0% 1 .66 1 0% .2 .33 9 -800% .25 .66 1 0 1 .66
uLeak 12 6 50% .14 .4 12 0% 0 0 8 33% .41 1 6 50% 1 .2 12 0% 0 0
AutoDispose 2 30 -1400% 0 0 1 50% 0 0 2 0% .01 .11 30 -1400% 0 0 24 -1100% 0 0
picasso 0 0 - .61 .39 0 0% .28 .8 0 - .68 .73 3 - .66 .02 0 - 0 0
skaffold-tools-for-java 1 1 0% 1 .71 1 0 0 0 1 0% .62 .71 2 -100% 0 0 1 0% 0 0
butterknife 4 4 0% 0 0 4 0% 0 0 0 100% 1 1 4 0% 0 0 4 0% 0 0
Totals 377 176 53% .31 .27 376 0% .57 .36 115 69% .39 .69 220 41% .76 .25 145 61% .91 .21
As an example of the the task that the models are being asked to do, consider the following line:
private static LeakInstanceTracker trackerSingleton = null;
A model should infer that this field is @Nullable, and such an annotation is present in the ground-truth human-annotated code; without an @Nullable annotation, NullAway complains that trackerSingleton is supposed to be non-null and yet it is assigned to null. Adding the annotation automatically would reduce the human effort to adapt this codebase to using NullAway, which is why we use warning reduction as a proxy for human effort.
7.2. RQ1
Our experiments compared the following approaches:
  • Graph Convolutional Networks (GCN) (Zhang et al., 2019a) is the common baseline model for node prediction tasks. We trained GCN models on the NaP-AST representation derived in section 5 (“NullGCN”).
  • The TDG encoding from prior work (Peng et al., 2022b) (“TDG GCN”).
  • A Graph Transformer Network (GTN) (Yun et al., 2019) is a model that is mathematically guaranteed to perform at least as well as a GCN, but extends its capabilities for graphs with multiple edge types. Our GTN model (“NullGTN”) is also trained on our NaP-AST encoding; the TDG encoding only has one edge type, so it gains no benefit from the GTN over the GCN.
  • Large Language Models (LLMs) like CodeQwen 1.5-7B-Chat and GPT-4 are a good baseline for coding tasks.
NullGCN
@@ 7.2. RQ1
On the NullAway benchmark, NullGCN yields 176 warnings vs 377 unannotated (53% reduction), with precision 0.31 / recall 0.27. While totals improve, the model regresses on some small projects (e.g., meal-planner 1 to 11, -1000%; AutoDispose 2 to 30, -1400%), indicating that misplaced annotations can introduce new warnings even as others are removed. While this is better than no inference at all, this result suggests to us that the hypergraph structure of the GTN (which does much better) is important to modeling where humans write annotations: the same graph is not as useful if the edges are not known to be of different types.
TDG GCN
TDG GCN achieves precision 0.57 / recall 0.39 but yields  0% warning reduction (376 vs 377). The inferences placed by this model were not accurate enough to reduce warnings on the projects. We hypothesize that the TDG chains are too disconnected to place annotations in a consistent manner. We speculate the TDG chains are better suited for use as an additional layer in the NaP-AST rather than an independent encoding. However, because the TDG encoding only has one edge type, it cannot benefit from the GTN model, and it is dramatically outperformed by the NaP-AST plus GTN combination. This shows that the dataflow hints in our name layer are critical for inferring pluggable types.
NullGTN
@@
We built our GTN model (NullGTN) by modifying a fork of the original code (Yun, 2023). It takes an average of about 16 hours to train the model. The largest input source code we evaluated (6,470 non-comment, non-blank lines) took about 5 minutes for inference. NullGTN achieved an F1 score of .95 on the test set during training. On the NullAway Benchmark, NullGTN discovered 69% of the human-written type qualifiers—most of them—with a precision of 39%. However, this only eliminated 69% of the NullAway warnings that a human would need to resolve to adapt these benchmarks to use NullAway—a large amount, but far from all of the warnings. One source of NullGTN’s inability to reduce warnings further is imprecision, which leads to extra annotations, which leads to extra warnings. Another cause of low warning reduction for NullGTN is interactions with libraries. NullGTN has no model of what libraries do, and unlike an LLM it cannot guess based on e.g., the names of library methods. The jib project (where NullGTN fails to remove many warnings despite good recall) uses libraries extensively. It is a modular system with many APIs and plugins, which the model was unaware of. Overall, these results are encouraging: NullGTN makes a good start on bringing these benchmarks into compliance with NullAway, and finds a majority (69%) of the annotations that the developers themselves keep in their codebases.
class LeakTracker {
private WeakReference<Object> gcIndicatorWeakReference;
LeakTracker(Class<?> eventClass, WeakReferenceGenerator g, String name, LeakType leakType) {
this.eventClass = eventClass;
this.weakReferenceGenerator = g;
this.name = name;
this.leakType = leakType;
Figure 7. Simplified example of code that CodeQwen fails to annotate correctly from uLeak’s LeakTracker.java.
GPT-4
To evaluate GPT-4’s suitability for pluggable type inference, we tested its performance on the same dataset used for our NaP-AST-based models. GPT-4 demonstrated moderate precision (76%) and recall (25%) compared to the graph-based models, indicating that while it can correctly annotate some types, it struggles to identify the majority of annotation sites. Despite slightly higher recall than CodeQwen, GPT-4’s more aggressive edits occasionally introduce warnings (e.g., picasso 0 to 3), which explains its lower net reduction.
GPT-4’s strength lies in its general coding abilities, but its lack of domain-specific understanding of pluggable type systems hinders its performance. Unlike NullGTN, GPT-4 does not incorporate graph-based dataflow hints, which are crucial for accurate type inference.
CodeQwen 1.5-7B-Chat
On the NullAway benchmark (Table 2), CodeQwen achieved 91% precision and 21% recall, reducing warnings by 61% overall. However, its edits can increase warnings on some projects (e.g., AutoDispose: 2 to 24 warnings), consistent with occasional unstable edits even at temperature 0.
The precision value is very high because it places relatively few annotations compared to the other models. This is because of CodeQwen’s greater conservatism in making changes. For example, in the code example in fig. 7, the gcIndicatorWeakReference field should be nullable because it is not initialized by the constructor. But since LLMs are trained to reason based on present tokens instead of absent ones, CodeQwen misses it.
Answer to RQ1
Across the 9-project benchmark (Table 2), NullGTN attains the highest recall (0.69) and warning reduction (69%) at moderate precision (0.39). CodeQwen has the highest precision (0.91) but lower recall (0.21) and reduces warnings by 61% overall. The other models and encodings all perform strictly worse than one of these two models. Graph models—especially NaP-AST + GTN—deliver the best overall downstream performance (highest recall and largest warning reduction). LLMs achieve very high precision (e.g., CodeQwen 91%) but lower recall. This suggests that a graph encoding is the right choice for type inference, which is in line with prior work on non-pluggable type inference (Peng et al., 2023b; Wei et al., 2023; Peng et al., 2023a, 2022a; Mir et al., 2022; Cui et al., 2021).
1020304050607080901000200200400400600600Training percentageWarnings
Figure 8. Warnings by percentage of data trained.
7.3. RQ2
We trained the GTN model with random samples of the dataset at 10% increments (10%, 20%, …) and ran the evaluation to see how much data is required for good results. We wanted to present the evaluation of the downstream task instead of presenting standard ML metrics (like F1 score) on their own, which can be potentially misleading: even a classifier that seems to perform well might not be useful for the downstream task.